OpenStudy (anonymous):

Write 32x5y5 with only one exponent. Use parentheses.

OpenStudy (anonymous):

32 = 2^5 ANSWER ( 2xy)^5

OpenStudy (anonymous):

(2xy)^5

OpenStudy (anonymous):

it is 32x^5y^5

OpenStudy (anonymous):

@doubledrive is right :)

OpenStudy (anonymous):

that is the question

OpenStudy (diyadiya):

\[\Large 32 = 2^5 \]\[\large 32x^5y^5 = (2)^5 (x)^5(y)^5 = (2xy)^5\]

OpenStudy (anonymous):

2Diyadiya i love ur pic

OpenStudy (diyadiya):

@doubledrive Haha Thanks

OpenStudy (diyadiya):

@chadandtricia_2009 32 = 2^5 because 32 = 2 x 2 x 2 x 2 x 2 = 2^5 Just incase if you don't know :)

OpenStudy (anonymous):

thank you